For a large sporting event the broadcasters sold

5757

ad slots for a total revenue of

​$148148

million. What was the mean price per ad​ slot?

The mean price per ad slot was

​$nothing

million.

​(Round to one decimal place as​ needed.)
Answer #1

Here n=57 and revenye=148 so the mean price =148/57=2.6

So tbe answrr is 2.6 is the price
Answer #2
Calculate mean price per slot Sale = 57 Total revenue = 148 Mean price = 148/57 Mean price = 2.6
answered by: Sehaj Kaur